# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ Abstract
In this review, we present our recent works on materials whose common point is the presence of electronic bands of very low dispersion, called “flat bands”, which are always the signature of an electronic confinement. A first part is devoted to the cases where this confinement is due to the long-range geometry of the defect-free structure. We have thus studied periodic approximant structures of quasiperiodic Penrose and octagonal tilings, and twisted bilayers of graphene or transition metal dichalcogenides (TMDs) whose rotation angle between the two layers assumes a special value, called “magic angle”. In these materials, the flat bands correspond to electronic states distributed over a very large number of atoms (several hundreds or even thousands of atoms) and are very sensitive to small structural distortions such as “heterostrain”. We have shown that their electronic transport properties cannot be described by usual Bloch-Boltzmann theories, because the interband terms of the velocity operator dominate the intraband terms as far as quantum diffusion is concerned. In the case of twisted bilayer graphene, flat bands can induce a magnetic state and other electron-electron correlation effects. The second part focuses on two-dimensional nanomaterials in the presence of local point defects that cause resonant electronic states (vacancies, adsorbed atoms or molecules). We present studies on monolayer graphene, twisted or Bernal bilayer graphene, carbon nanotubes, monolayer and multilayer black phosphorene, and monolayer TMDs. A recent result is the discovery that the selective functionalization of a Bernal bilayer graphene sublattice leads to a metallic or insulating behavior depending on the functionalized sublattice type. This result, which seems to be confirmed by very recent experimental measurements, suggests that functionalization can be a key parameter to control the electronic properties of two-dimensional materials.
@@ 1 Introduction
For around two decades, the physics of 2D materials has been steadily progressing, both in terms of the fundamental insights it provides and its potential applications [1, 2, 3, 4]. As far as their electronic properties are concerned, a central aspect is the existence of electronic states with very low energy dispersion, known as “flat-band states” or “almost flat-band states”. The origin of these states is diverse. On the one hand, they can be caused by surfaces, edges or local defects in the crystallographic structure, such as adsorbates. On the other hand, they may also be the consequence of electronic confinement mechanisms by complex geometries, such as in moiré patterns in the twisted bilayers 2D materials. Another important class of materials with flat bands which have been much studied both theoretically and experimentally consists of quasicrystals and their so-called crystalline approximants. In their perfect structure these systems generate flat bands at all energies that are related to the existence of critical states. These materials provide an important example for systems with flat bands, diffferent from, but connected to moiré patterns in 2D materials. In this review, we present some of the flat-band states of these two types and discuss aspects of their electronic properties that we consider essential and we have been working on. Let us emphasize that our approach uses realistic tight-binding Hamiltonian models so that we can treat aspects which are relevant to actual materials.
@@
We note that beyond the 2D materials and quasicrystals which are considered in this review there are many other partially overlapping notions of flat bands in condensed matter physics, see, e.g., Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], including flat-band ferromagnetism in the Hubbard model [5] and localized eigenstates arising from flat bands in quantum spin systems [6, 7] and photonic lattices [8, 9]. As discussed in the literature [19, 20, 13, 15, 21, 14, 16, 17, 18, 22, 23, 24, 25], these flat-band states may result from topological properties of these materials. We will not enter into any further details about these systems here, but instead refer to the aforementioned review articles.
@@
Figure 1: Schematic transport regime diagram versus the elastic or inelastic mean-free path $L$ due to defects. $a$ is the interatomic distance and $L_{w}$ the extent of wave packet of charge carriers.
@@
Flat bands are associated to confinement which is expected to occur on some typical length scale. This length should be important for electronic structure (in the paramagnetic or magnetic state) and also for transport. The usual theories to understand the electronic transport properties of materials for which electronic interactions are not strong, such as metals or semiconductors, are based on the semi-classical approximation and Bloch-Boltzmann theories [26]. These approaches use the concept of wave packets which travel at a given velocity (equal to the band velocity) between two scattering events. As we show in this review, the minimal extent $L_{w}$ of a wave packet depends on the properties of the perfect structure and may play a prominent role, especially when the band velocity becomes small. The semi-classical model applies well provided that the spatial extent $L_{w}$ of charge carrier wave packets is small compared to the mean free path $L$, i.e., the average distance between two collisions on static scatterers (point defects) or dynamic scatterers (electron/phonon interactions or effect of an external magnetic field). In a crystal, electronic energy bands $E_{n}(\mathbf{k})$ can be defined and the semi-classical condition is thus written,
@@
where $V_{\mathrm{B}}$ is the Boltzmann (intraband) velocity in the $i$-direction and $\tau$ the elastic or inelastic scattering time. As we will see in the following, $L_{w}$ results from a complex effect of the structure. It depends on $E_{F}$ but may also depend on $\tau$. When specific mechanisms localize or confine charge carriers, the velocity $V_{\mathrm{B}}$ can be greatly reduced. And then, if the localization mechanism confines the states to large sizes, typically much larger than the interatomic distance, $L_{w}$ is large and the condition (1) breaks down, rendering the semi-classical approximation invalid (Fig. 1). This unusual localization mechanism, called “Small Velocity Regime” (SVR), has been highlighted in the approximants of aluminum-based quasicrystals to qualiltatively understand their anomalous conductivity, which, while being gap-free, behave like semiconductors [27, 28]. More recently, this regime has been also demonstrated in the electronic states confined by magic-angle twisted bilayer graphene (TBG) [29]. SVR, which is not due to defects in the crystallographic structure, is the first source of the electronic flat bands presented here. The second origin, which is better known but also very interesting, both from a theoretical point of view and for its potential applications, is the existence of localized states due to local defects, in particular resonant defects such as atoms or molecules adsorbed onto the surface of the 2D material (functionalization).
@@
The paper is organized as follows. The section 2 presents cases where the flat bands are due to the particular geometry of the systems without defects: quasiperiodic tilings (Sec. 2.2), twisted bilayer graphene (Sec. 2.3) and twisted bilayer transition metal dichalcogenides (TMDs) (Sec. 2.4). Section 3 focuses on quantum transport studies in systems where flat bands arise from static defects: monolayer and bilayer graphene (Bernal AB bilayer and twisted Bilayer Graphene (TBG)) with defects randomly distributed throughout the structure (Sec. 3.2) or selectively in specific atomic sublattices (Sec. 3.3), metallic Single Wall Carbon Nanotubes (SWCNTs) (Sec. 3.4), monolayer TMDs (Sec. 3.5.1) and monolayer and multilayer black phosphorene (Sec. 3.5.2). Our aim is to describe the properties of real materials as much as possible. Therefore, we pay particular attention to the tight-binding (TB) model of the studied systems. In the framework of the linear response theory and the Kubo-Greenwood approach [30, 31, 32], to study quantum diffusion, we use numerical methods in reciprocal space for systems without defects and whose unit cell is not too large, and a real-space method for very large systems. These numerical methods to calculate quantum transport by recursion [33, 34, 35, 36, 37, 38] or by diagonalization [27, 29] are presented in the A. These methods include all quantum effects for non-interacting charge carriers, including multiple scattering effects.
@@
Immediately after the discovery by D. Shechtman, I. Blech, D. Gratias, and J. W. Cahn [39] of quasiperiodic intermetallics, one major question was raised about the physical properties of phases with this new type of order, non-periodic but with long-range order. In particular, one expected that electronic and thermal properties of quasicrystals could be deeply affected [40, 41]. The description of electrons or phonons in periodic phases rests on the Bloch theorem which cannot be applied to a quasiperiodic structure. Within several decades a series of new quasiperiodic phases, periodic approximants and related complex intermetallic alloys were discovered and intensively studied. These investigations have taught us that indeed electron and phonon properties could be strongly affected by this new type of order with respect to usual intermetallic alloys (crystals or amorphous alloys).
The first quasiperiodic alloys, found in Al-Mn systems, were metastable and contained many structural defects [39]. As a consequence they had conduction properties similar to those of amorphous metals with resistivities in the range 100–500 $\mu\Omega$cm. In 1986, the first stable icosahedral phase was discovered in AlLiCu systems. This phase was still defective and although its resistivity was higher (800 $\rm\mu\Omega$cm), it was still comparable to that of amorphous metals. The real breakthrough came with the discovery of the stable AlCuFe icosahedral phase, having a high structural order. The resistivity of these very well ordered systems were very high, of the order of $\rm 10^{4}\,{\rm\mu\Omega cm}$ [42, 43], which gave a considerable interest in their conduction properties. Within a few years several important electronic characteristics of these phases were experimentally demonstrated. The density of states in AlCuFe was smaller than in Al, about one third of that of pure Al, but still largely metallic. The conductivity presented a set of characteristics that were either those of semiconductors or those of normal metals. In particular weak-localization effects were observed which are typical of amorphous metals. Yet the conductivity increases with the number of defects just as in semiconductors. Optical measurements showed that the Drude peak, characteristic of normal metals, was absent. In 1993 another breakthrough was the discovery of AlPdRe which had resistivities in the range of $\rm 10^{6}\,{\rm\mu\Omega cm}$ [44, 45, 46]. This system gave the possibility of studying a metal-insulator transition in a quasiperiodic phase. There are still many questions concerning electronic transport in AlPdRe phases [47].
Since the discovery of quasicrystals, the view of the role of quasiperiodic order has evolved. On one hand, the long-range quasiperiodic order can induce electronic states which are neither localized nor extended and called “critical states” (see Refs. [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 35, 36, 58, 37, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and Refs. therein). On the other hand, for the electronic properties of most known alloys, it appears that the medium-range order, spanning one or a few nanometers, is the actual length scale that determines these properties. This observation has led the scientific community to adopt a larger point of view and consider quasicrystals as an example of a larger class. This class of complex metallic alloys contains quasicrystals, approximants and alloys with large and complex unit cells with possibly hundreds of atoms in the unit cell.
@@
Graphene is a 2D carbon material which takes the form of a planar lattice of $sp^{2}$ bonded atoms. Since it could be isolated in 2004 [69, 70, 71, 72, 73], it is of great interest both theoretically and experimentally. Its honeycomb lattice consists of two sublattices which give a specific property to the wave function, the so-called chirality. The linear dispersion relation close to the charge neutrality point implies that electron states are well described by an effective massless Dirac equation at low energy (Dirac cone) [74, 72, 73, 75, 76]. The properties of electrons in graphene, obtained from the corresponding Dirac equation, are fundamentally different from those deriving from the Schrödinger equation. In particular, the quantum Hall effect is quantized in integer and half-integer multiples of the fundamental conductance quantum [72, 73], and can even be observed at room temperature [77]. Another example of the unique behavior of Dirac electrons is the so-called Klein paradox which is intimately related to the chirality of their wavefunction [78]. The Klein paradox is the fact that, in a one-dimensional configuration, a potential barrier is perfectly transparent for electrons. As a result, it is challenging to localize Dirac electrons using an electrostatic potential. However, achieving such confinement could be highly beneficial, particularly for the development of elementary devices.
@@
Graphene is a zero gap material, which may limit its potential applications. However, electronic transport in graphene is sensitive to static defects such as frozen ripples, screened charged impurities, or local defects like vacancies or adsorbates [79, 80, 81, 82, 83, 84, 85, 86, 38]. One way to create and tune a gap in graphene is the selective functionalization which has been used, e.g., with hydrogen adsorption on a moiré pattern of Graphene-Ir(111) [87].
@@
The study of graphene has also opened up the way to the study of the broader families of 2D materials such as hexagonal boron nitrite (h-BN), transition metal dichalcogenide (TMDs), phosphorene, thin oxides… In fact it opens up the way to “flatland” [88], a new paradigm for electronics and optoelectronics. 2D materials emerge as the cutting edge of physical and chemical science and engineering. Stacking layered materials is also a very powerful method to tailor their electronic properties [1]. The properties not only depend on the choice of materials to be stacked but also on the details of the relative arrangement of the layers, e.g., the rotation angle $\theta$ between the two layers in twisted bilayers of graphene (TBG) [89].It has thus been shown theoretically [89, 90, 91, 92, 93, 94, 95] and experimentally [96, 97] that TBGs, forming a moiré pattern [98, 99, 100, 101] at “magic angles” [91] confine electrons in a tunable way in terms of the angle of rotation $\theta$ of one layer with respect to the other. In 2018, it has been experimentally proven that this localization by geometry induces strong electronic correlations and a superconducting state for specific angles [102, 103]. The rotation angle $\theta$ is a key parameter to fix the geometry and the properties of the flat bands due to the moiré pattern, but it has also been shown that other effects are very important in particular, the “heterostrain” (axial strain in one of the two layers) [104, 105, 106] and the atomic relaxation [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122]. Despite numerous studies on these new systems (see, e.g., Refs. [29, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133]), the consequences of this electronic localization by a moiré pattern on electrical transport properties are still poorly known. Moreover in 2018, “quasiperiodic Dirac states” have been observed by diffraction in bilayer graphene with a rotation angle $\theta=30^{\circ}$ [134], thus showing the great richness and variety of possible phases in these systems. For an overview of the electronic properties of TBGs and their potential applications, see, e.g., Refs. [135, 136, 137, 138, 139, 140, 141, 142, 14]. Since the discovery of TBG, many other homo- or hetero-twisted 2D materials have been discovered, in particular based on semiconductor materials such as TMDs and h-BN  [143, 144, 145, 146], In the latter, depending on the rotation angle, flat bands and quasiperiodic states may also exist [147, 148, 149, 150] (see also Refs. in Sec. 2.4).

### v2

@@ Abstract
In this review, we present recent works on materials whose common point is the presence of electronic bands of very low dispersion, called “flat bands”, which are due to specific atomic order effects without electron interactions. These states are always indicative of some form of confinement and have significant consequences on the electronic structure, transport properties and magnetism of these materials. A first part is devoted to the cases where this confinement is due to the long-range geometry of the defect-free structure. We have thus studied periodic approximant structures of quasiperiodic Penrose and octagonal tilings, and twisted bilayers of graphene or transition metal dichalcogenides (TMDs) whose rotation angle between the two layers assumes a special value, called “magic angle”. In these materials, the flat bands correspond to electronic states distributed over a very large number of atoms (several hundreds or even thousands of atoms) and are very sensitive to small structural distortions such as “heterostrain”. We have shown that their electronic transport properties cannot be described by usual Bloch-Boltzmann theories, because the interband terms of the velocity operator dominate the intraband terms as far as quantum diffusion is concerned. In the case of twisted bilayer graphene, flat bands can induce a magnetic state and other electron-electron correlation effects. The second part focuses on two-dimensional nanomaterials in the presence of local point defects that cause resonant electronic states (vacancies, adsorbed atoms or molecules). We present studies on monolayer graphene, twisted or Bernal bilayer graphene, carbon nanotubes, monolayer and multilayer black phosphorene, and monolayer TMDs. A recent result is the discovery that the selective functionalization of a Bernal bilayer graphene sublattice leads to a metallic or insulating behavior depending on the functionalized sublattice type. This result, which seems to be confirmed by very recent experimental measurements, suggests that functionalization can be a key parameter to control the electronic properties of two-dimensional materials.
@@ 1 Introduction
For around two decades, the physics of 2D materials has been steadily progressing, both in terms of the fundamental insights it provides and its potential applications Geim13 , Lin2023 , Lei2022 , KUMBHAKAR2023106671 . As far as their electronic properties are concerned, a central aspect is the existence of electronic states with very low energy dispersion, known as “flat-band states” or “almost flat-band states”. The origin of these states is diverse. On the one hand, they can be caused by surfaces, edges or local defects in the crystallographic structure, such as adsorbates. On the other hand, they may also be the consequence of electronic confinement mechanisms by complex geometries, such as in moiré patterns in the twisted bilayer 2D materials. Another important class of materials with flat bands which have been much studied both theoretically and experimentally consists of quasicrystals and their so-called crystalline approximants. In their perfect structure these systems generate flat bands at all energies that are related to the existence of critical states. These materials provide an important example for systems with flat bands, diffferent from, but connected to moiré patterns in 2D materials. In this review, we present some of the flat-band states of these two types and discuss aspects of their electronic properties that we consider essential and we have been working on. Let us emphasize that our approach uses realistic tight-binding Hamiltonian models so that we can treat aspects which are relevant to actual materials.
@@
We note that beyond the 2D materials and quasicrystals which are considered in this review there are many other partially overlapping notions of flat bands in condensed matter physics, see, e.g., Refs. Tasaki1998 , Richter2004 , Derzhko2007 , Leykam18 , DanieliAndreanovLeykamFlach2024 , Liu20 , Liu22 , Mitscherling22 , Regnault2022 , Bernevig24 , Calugaru22 , Thumin23 , Thumin25 , Jiang25 , including flat-band ferromagnetism in the Hubbard model Tasaki1998 and localized eigenstates arising from flat bands in quantum spin systems Richter2004 , Derzhko2007 and photonic lattices Leykam18 , DanieliAndreanovLeykamFlach2024 . As discussed in the literature Heikkil16 , Song19 , Liu21_topo , Regnault2022 , Calugaru22 , Song22 , Bernevig24 , Thumin23 , Thumin25 , NunezRegueiro24 , Jiang25 , Kwan25 , Iugov25 , Thompson25 , Mesple25 , Duncan20 , Manna24 , these flat-band states and the exotic properties such as superconductivity may result from topological effects. We will not enter into any further details about these systems here, but instead refer to the aforementioned review articles.
@@
Figure 1: Schematic transport regime diagram versus the elastic or inelastic mean free path $L$ due to defects. $a$ is the interatomic distance and $L_{w}$ the extent of wave packets of charge carriers.
@@
Flat bands are associated to confinement which is expected to occur on some typical length scale. This length should be important for electronic structure (in the paramagnetic or magnetic state) and also for transport. The usual theories to understand the electronic transport properties of materials for which electronic interactions are not strong, such as metals or semiconductors, are based on the semi-classical approximation and Bloch-Boltzmann theories Ashcroft76 . These approaches use the concept of wave packets which travel at a given velocity (equal to the band velocity) between two scattering events. As we show in this review, the minimal extent $L_{w}$ of a wave packet depends on the properties of the perfect structure and may play a prominent role, especially when the band velocity becomes small. The semi-classical model applies well provided that the spatial extent $L_{w}$ of charge carrier wave packets is small compared to the mean free path $L$, i.e., the average distance between two collisions on static scatterers (point defects) or dynamic scatterers (electron/phonon interactions or effect of an external magnetic field). In a crystal, electronic energy bands $E_{n}(\mathbf{k})$ can be defined and the semi-classical condition is thus written,
@@
where $V_{\mathrm{B}}$ is the Boltzmann (intraband) velocity in the $i$-direction and $\tau$ the elastic or inelastic scattering time. As we will see in the following, $L_{w}$ results from a complex effect of the structure. It depends on $E_{F}$ but may also depend on $\tau$. When specific mechanisms localize or confine charge carriers, the velocity $V_{\mathrm{B}}$ can be greatly reduced. If the localization mechanism confines the states to large sizes, typically much larger than the interatomic distance, then $L_{w}$ is large and the condition (1) breaks down, rendering the semi-classical approximation invalid (Fig. 1). This unusual localization mechanism, called “Small Velocity Regime” (SVR), has been highlighted in the approximants of aluminum-based quasicrystals to qualiltatively understand their anomalous conductivity, which, while being gap-free, behave like semiconductors Trambly06 , Mayou08 . More recently, this regime has been also demonstrated in the electronic states confined by magic-angle twisted bilayer graphene (TBG) Trambly16 . SVR, which is not due to defects in the crystallographic structure, is the first source of the electronic flat bands presented here. The second origin, which is better known but also very interesting, both from a theoretical point of view and for its potential applications, is the existence of localized states due to local defects, in particular resonant defects such as atoms or molecules adsorbed onto the surface of the 2D material (functionalization).
@@
The paper is organized as follows. The section 2 presents cases where the flat bands are due to the particular geometry of the systems without defects: quasiperiodic tilings (Sec. 2.2), twisted bilayer graphene (Sec. 2.3) and twisted bilayer transition metal dichalcogenides (TMDs) (Sec. 2.4). Section 3 focuses on quantum transport studies in systems where flat bands arise from static defects: monolayer and bilayer graphene (Bernal AB bilayer and twisted Bilayer Graphene (TBG)) with defects randomly distributed throughout the structure (Sec. 3.2) or selectively in specific atomic sublattices (Sec. 3.3), metallic Single Wall Carbon Nanotubes (SWCNTs) (Sec. 3.4), monolayer TMDs (Sec. 3.5.1) and monolayer and multilayer black phosphorene (Sec. 3.5.2). Our aim is to describe the properties of real materials as much as possible. Therefore, we pay particular attention to the tight-binding (TB) model of the studied systems. In the framework of the linear response theory and the Kubo-Greenwood approach Kubo57 , Greenwood58 , Kubo91 , to study quantum diffusion, we use numerical methods in reciprocal space for systems without defects and whose unit cell is not too large, and a real-space method for very large systems. These numerical methods to calculate quantum transport by recursion Mayou88 , Mayou95 , Roche97 , Roche99 , Triozon02 , Trambly13 or by diagonalization Trambly06 , Trambly16 are presented in the A. These methods include all quantum effects for non-interacting charge carriers, including multiple scattering effects.
@@
Immediately after the discovery by D. Shechtman, I. Blech, D. Gratias, and J. W. Cahn shechtman84 of quasiperiodic intermetallics, one major question was raised about the physical properties of phases with this new type of order, non-periodic but with long-range order. In particular, one expected that electronic and thermal properties of quasicrystals could be deeply affected Poon92 , Berger94 . The description of electrons or phonons in periodic phases rests on the Bloch theorem which cannot be applied to a quasiperiodic structure. Within several decades a series of new quasiperiodic phases, periodic approximants and related complex intermetallic alloys were discovered and intensively studied. These investigations have taught us that indeed electron and phonon properties could be strongly affected by this new type of order with respect to usual intermetallic alloys (crystals or amorphous alloys).
[…]
Since the discovery of quasicrystals, the view of the role of quasiperiodic order has evolved. On one hand, the long-range quasiperiodic order can induce electronic states which are neither localized nor extended and called “critical states” (see Refs. Kohmoto86 , Fujiwara89_Fibo , Sire90 , Passaro92 , Sire94 , Mayou94 , Guarneri94 , Piechon96 , SchulzBaldes98 , Zhong98 , Roche97 , Roche99 , Mayou00 , Triozon02 , Jagannathan07 , Trambly11_QC , Trambly14_QC , Trambly17 , Collins17 , Chiaracane21 , Grimm21 , Jagannathan23 , Jagannathan24 , Bellissard24 and Refs. therein). On the other hand, for the electronic properties of most known alloys, it appears that the medium-range order, spanning one or a few nanometers, is the actual length scale that determines these properties. This observation has led the scientific community to adopt a larger point of view and consider quasicrystals as an example of a larger class. This class of complex metallic alloys contains quasicrystals, approximants and alloys with large and complex unit cells with possibly hundreds of atoms in the unit cell.
@@
Graphene is a 2D carbon material which takes the form of a planar lattice of $sp^{2}$ bonded atoms. Since it could be isolated in 2004 Novoselov04 , Berger04 , Hashimoto04 , Novoselov2005 , Zhang05 , it is of great interest both theoretically and experimentally. Its honeycomb lattice consists of two sublattices which give a specific property to the wave function, the so-called chirality. The linear dispersion relation close to the charge neutrality point implies that electron states are well described by an effective massless Dirac equation at low energy (Dirac cone) Wallace47 , Novoselov2005 , Zhang05 , Berger06 , Sadowski06 . The properties of electrons in graphene, obtained from the corresponding Dirac equation, are fundamentally different from those deriving from the Schrödinger equation. In particular, the quantum Hall effect is quantized in integer and half-integer multiples of the fundamental conductance quantum Novoselov2005 , Zhang05 , and can even be observed at room temperature Novoselov2007 . Another example of the unique behavior of Dirac electrons is the so-called Klein paradox which is intimately related to the chirality of their wavefunction Katsnelson2006 . The Klein paradox is the fact that, in a one-dimensional configuration, a potential barrier is perfectly transparent for electrons. As a result, it is challenging to localize Dirac electrons using an electrostatic potential. However, achieving such confinement could be highly beneficial, particularly for the development of elementary devices.
@@
Graphene is a zero gap material, which may limit its potential applications. However, electronic transport in graphene is sensitive to static defects such as frozen ripples, screened charged impurities, or local defects like vacancies or adsorbates Peres06 , Pereira06 , Pereira08a , Lherbier08 , Lherbier08b , Leconte10 , Leconte11 , Skrypnyk10 , Trambly13 . One way to create and tune a gap in graphene is the selective functionalization which has been used, e.g., with hydrogen adsorption on a moiré pattern of Graphene-Ir(111) Jorgensen16 .
@@
The study of graphene has also opened up the way to the study of the broader families of 2D materials such as hexagonal boron nitrite (h-BN), transition metal dichalcogenide (TMDs), phosphorene, thin oxides… In fact it opens up the way to “flatland” Novoselov11 , a new paradigm for electronics and optoelectronics. 2D materials emerge as the cutting edge of physical and chemical science and engineering. Stacking layered materials is also a very powerful method to tailor their electronic properties Geim13 . The properties not only depend on the choice of materials to be stacked but also on the details of the relative arrangement of the layers, e.g., the rotation angle $\theta$ between the two layers in twisted bilayers of graphene (TBG) LopesdosSantos07 .It has thus been shown theoretically LopesdosSantos07 , Trambly10 , Bistritzer10 , SuarezMorell10 , Bistritzer11 , Trambly12 , Santos12 and experimentally Li10 , Brihuega12 that TBGs, forming a moiré pattern Campanera07 , Mele10 , Gratias23 , Gratias20 at “magic angles” Bistritzer10 confine electrons in a tunable way in terms of the angle of rotation $\theta$ of one layer with respect to the other. In 2018, it has been experimentally proven that this localization by geometry induces strong electronic correlations and a superconducting state for specific angles Cao18a , Cao18b , Choi2019 , Jiang2019 , Xie19 . The rotation angle $\theta$ is a key parameter to fix the geometry and the properties of the flat bands due to the moiré pattern, but it has also been shown that other effects are very important in particular, the “heterostrain” (axial strain in one of the two layers) Huder18 , Bi19 , Mesple21 (see also Kerelsky2019 , Gao21 , Pantaleon21 , Mannai21 , Parker21 , XIONG2023129048 , Escudero24 , Yu2025 ) and the atomic relaxation Uchida14 , Wijk15 , Dai16 , Nam17 , Gargiulo17 , Choi18 , Angeli18 , Lin18 , Lin18b , Lin18c , Koshino18 , Lucignano19 , Yoo19 , Guinea19 , Liu21 , Nguyen21 , Yananose21 , Kazmierczak21 , Leconte22 , Ezzi24 , Ceferino24 . Despite numerous studies on these new systems (see, e.g., Refs. Trambly16 , Chung18 , Andelkovic18 , Omid20 , Ciepielewski22 , Ciepielewski24 , Sinha24 , guerrero24_tBLG_disorder_Kubo , Rai24 , Kim24 , GobboKhun25 , Guerrero25 ), the consequences of this electronic localization by a moiré pattern on electrical transport properties are still poorly known. Moreover in 2018, “quasiperiodic Dirac states” have been observed by diffraction in bilayer graphene with a rotation angle $\theta=30^{\circ}$ Ahn18 , Yao18 , Yu19 , Yan19 , thus showing the great richness and variety of possible phases in these systems. For an overview of the electronic properties of TBGs and their potential applications, see, e.g., Refs. Rozhkov16 , Catarina19 , Wang19 , andrei2020graphene , Carr20 , Mogera20 , Ledwith21 , Sun24_revue , Bernevig24 . Since the discovery of TBG, many other homo- or hetero-twisted 2D materials have been discovered, in particular based on semiconductor materials such as TMDs and h-BN  Wang15_reviewTMD , Liu15_reviewTMD , Duong17_review , Latil23 , In the latter, depending on the rotation angle, flat bands and quasiperiodic states may also exist Xian19 , Ochoa20 , Walet21 , Sponza24 (see also Refs. in Sec. 2.4).